OpenStudy (anonymous):

find the value of 3/(sqrt)8

OpenStudy (mimi_x3):

\[\frac{3*\sqrt{8}}{\sqrt{8}*\sqrt{8}} = \frac{3\sqrt{8}}{8} =>\frac{3}{8} \sqrt{8} \]